# RUSplatting: Robust 3D Gaussian Splatting for Sparse-View Underwater Scene Reconstruction

### v1

@@ Abstract
Reconstructing high-fidelity underwater scenes remains a challenging task due to light absorption, scattering, and limited visibility inherent in aquatic environments. This paper presents an enhanced Gaussian Splatting-based framework that improves both the visual quality and geometric accuracy of deep underwater rendering. We propose decoupled learning for RGB channels, guided by the physics of underwater attenuation, to enable more accurate colour restoration. To address sparse-view limitations and improve view consistency, we introduce a frame interpolation strategy with a novel adaptive weighting scheme. Additionally, we introduce a new loss function aimed at reducing noise while preserving edges, which is essential for deep-sea content. We also release a newly collected dataset, Submerged3D, captured specifically in deep-sea environments. Experimental results demonstrate that our framework consistently outperforms state-of-the-art methods with PSNR gains up to 1.90dB, delivering superior perceptual quality and robustness, and offering promising directions for marine robotics and underwater visual analytics.
@@ 1 Introduction
Advancements in Novel View Synthesis (NVS) have introduced new possibilities for reconstructing and visualising environments in 3D space from multi-view image data. These methods allow the generation of unseen viewpoints using limited training images. A notable breakthrough in NVS is Neural Radiance Fields (NeRF) [21], which implicitly models 3D scenes. Despite their impressive performance, NeRFs require substantial training time and are impractical for real-time or interactive applications [2, 28]. Moreover, their representations are inherently difficult to interpret or modify intuitively [28], limiting their flexibility in dynamic environments, such as underwater. In contrast, 3D Gaussian Splatting (3DGS) [15] reconstructs scenes explicitly using Gaussian point sets, offering faster training, higher rendering speeds and refresh rates, as well as more flexible scene manipulation.
@@
Nevertheless, most existing NeRF-based and 3DGS-based methods focus on reconstructing scenes in clear media and therefore struggle to produce high-quality results in hazy and distorted environments like underwater. Due to the physical properties of water, underwater imaging is affected by attenuation and scattering, where the former causes colour casting and the latter results in low contrast [20]. It is therefore essential to account for these properties when modelling the entire scene.
@@
In the face of these challenges, we propose a new framework: Robust Underwater scene representation using 3D Gaussian Splatting (RUSplatting), which demonstrates greater robustness than previous works, such as UW-GS [27] and WaterSplatting [18]. Previous approaches integrate medium effects via affine transformations of Gaussian attributes but assume dense, temporally coherent input frames, limiting effectiveness under sparse-view conditions. They also do not address distortions common in deep-sea imaging, such as strong noise and severe colour imbalance.
@@
Our proposed RUSplatting framework, shown in Fig. 1, further advances underwater 3D reconstruction by explicitly decoupling the estimation of attenuation and scattering parameters for individual RGB channels instead of treating them uniformly, enabling more accurate and realistic colour restoration. As discussed in [23], the light attenuation underwater has different degrees of changes due to the wavelength. Red light, which has a longer wavelength, is absorbed more rapidly than green and blue light. Furthermore, to mitigate image noise prevalent in deep-sea conditions, we introduce an Edge-Aware Smoothness Loss (ESL) to directly optimise the reconstruction pipeline, dynamically reducing noise while preserving critical geometric edges. Additionally, to address the inherent challenges of sparse-view datasets, we incorporate an Intermediate Frame Interpolation (IFI) mechanism, enriching viewpoint coverage and increasing initialised Gaussian points. However, since interpolated frames may contain artefacts that could deteriorate the learning process, we propose an Adaptive Frame Weighting (AFW) strategy that adjusts the loss based on the predictive uncertainty [13, 14]. Finally, we introduce a new dataset, Submerged3D, capturing detailed objects in deep-sea environments, supporting future research in underwater robotics, archaeology, and marine biology. In summary, our key contributions are as follows:
@@
Extensive experiments on both existing and new datasets demonstrate that RUSplatting consistently outperforms state-of-the-art methods across various evaluation metrics, with PSNR improvements of up to 21.37%, 4.99% and 5.09% compared to SeeThru-NeRF [17], WaterSplatting [18] and UW-GS [27], respectively.
@@ 2 Related Work
Processing underwater images remains a major challenge in underwater computer vision due to colour distortion and noise arising from water’s physical properties and its impact on light propagation [7]. The Underwater Image Formation Model (UIFM) [3, 4] was introduced to describe this process, accounting for light absorption and backscattering. While the original UIFM assumes a homogeneous water column, the Revised UIFM (R-UIFM) [1] improves realism by allowing spatial variability in water properties, making it widely adopted in underwater image processing. For example, PRWNet [12] uses a deep learning-based multi-stage refinement approach that enhances image quality by addressing spatial and frequency domain degradations through Wavelet Boost Learning. However, the survey in [10] shows that applying enhancement prior to 3D reconstruction results in poorer performance compared to embedding the physical properties of water directly into the reconstruction process.
@@
While the NeRF-based and GS-based technologies have rapidly advanced, most struggle in scattering media such as fog or water. To tackle this limitation, ScatterNeRF [24] was introduced to effectively separate objects from the scattering medium in foggy environments. SeaThru-NeRF [17] extends this line of research by estimating water medium coefficients using a novel architecture grounded in the Revised Underwater Image Formation Model (R-UIFM) [1]. AquaNeRF [8] improves SeaThru-NeRF with single surface rendering to avoid distractors.  Tang et al. [26] propose a neural representation that incorporates dynamic rendering and tone mapping to enable more accurate reconstruction of underwater scenes under varying illumination conditions.
For 3DGS-based methods, UW-GS [27] and WaterSplatting [18] employ MLP-based colour appearance modules to address colour casting. In contrast, SeaSplat [29] optimises backscatter parameters using a dark channel prior loss. UW-GS [27], UDR-GS [5], and Aquatic-GS [20] leverage DepthAnything [30] to estimate depth from 2D inputs and incorporate it into their loss functions. UW-GS [27] further improves reconstruction of distant objects through a physics-driven density control module.
@@ 3 Preliminaries
The 3DGS [15] employs a collection of spatial Gaussian cloud points to represent a 3D scene. To synthesise a novel view, these Gaussians are projected onto a 2D image plane. The projected Gaussian is defined as
@@
Compared to clear, onshore scenes, where scattering and absorption caused by the medium (i.e., air) can often be neglected, underwater scenes are heavily affected by these parameters, since they significantly degrade reconstruction quality [26]. Therefore, it is crucial to consider these medium parameters during reconstruction. As modelled in [1], the intensity of the underwater image is defined as follows:
@@
Here $\beta^{d}$ and $\beta^{b}$ are two colour medium coefficients, and $z$ is the distance.
@@ 4 Methodology
The detailed structure of the proposed framework is presented in Fig. 1. The model estimates underwater medium parameters using an MLP with encoded depth and viewpoint position information as inputs. Accordingly, the colours of the spatial 3D Gaussians are modified through the affine transformation. Then, the modified Gaussians, including additional Gaussians introduced by the frame interpolation process with learnable uncertainty weights, are projected onto the 2D plane. During training, an edge-aware smoothness loss is integrated into the model as a part of the loss function to remove noise from underwater images, and in the meantime maintaining geometric details.
@@ 4.1 Colour-Channel Decoupling
As discussed in numerous underwater studies [20, 19, 10], the attenuation rates differ across red, green, and blue light channels. Therefore, the medium parameters for the RGB channels cannot be treated identically. To overcome this and achieve more accurate colour reconstruction, we decouple the channel-wise learning of the underwater medium. Consequently, the attenuation ($T^{D}$) and backscatter ($T^{B}$) factors for each colour channel $ch\in\{r,g,b\}$ are defined as:
@@
Previous works use a 2-layer MLP [16] to estimate the complex underwater medium parameters, however, it may fall short in capturing such complicated patterns as the study indicated in [9]. Thus, we employ a deeper MLP with 5 layers, selected based on a hyperparameter sensitivity analysis (see supplementary material for details). The final corrected colour of Gaussian $i$, $c_{i,ch}^{m}(v)$, is obtained through an affine transformation, as presented below, where $c_{i,ch}(v)$ is the observed view-dependent colour.
@@ 4.2 Intermediate Frame Interpolation (IFI)
To address the limitations introduced by sparse views, we adopt a frame interpolation technique to synthesise intermediate frames between adjacent input images, leveraging them in two key aspects: i) enhancing the initial point cloud and ii) improving the RUSplatting optimisation, where we propose Adaptive Frame Weighting (AFW). Although various methods are available and should not significantly affect the final 3DGS results, we use Real-time Intermediate Flow Estimation (RIFE) [11], a state-of-the-art deep learning-based approach that employs bidirectional optical flow and knowledge distillation to enable real-time processing.
@@
The initial point cloud acts as a rough geometric framework for the final model. Its quality and density directly influence the accuracy of the rendering. By incorporating interpolated frames, additional initialisation points are introduced, enabling a denser distribution of Gaussians and improving the overall reconstruction precision and detail. Without IFI, the model accesses an average of 17,635 initialisation points across all datasets. After adopting IFI, this number increases significantly by 55.7% to an average of 27,461 points. Detailed results for each dataset are provided in the supplementary material.
@@
In sparse view scenarios, adjacent frames are significantly spaced apart, making frame interpolation challenging. The large spatial shifts between frames may introduce artefacts in certain geometric structures of the interpolated frames, which may degrade the quality of 3D reconstruction during optimisation when these frames are used. Inspired by [13, 14], we propose an AFW mechanism, introducing an uncertainty learnable parameter $\gamma$ for each interpolated frame $f$ that is optimised along with the training. Accordingly, the loss function for the interpolated frame $L_{f}$ is updated as follows.
@@ 4.3 Edge-Aware Smoothness Loss (ESL)
where $(i,j)$ indexes the spatial coordinates in the image plane, $I$ denotes the input image and the $D$ denotes the pseudo depth map generated from the Depth-Anything-V2 [31]. The edge-aware weights $w^{x}_{i,j}$ and $w^{y}_{i,j}$, corresponding to the horizontal and vertical directions, respectively, are computed using depth discontinuities,
@@ 4.4 Loss Function
To evaluate the differences between the reconstructed image and the ground truth image, the $L_{\text{Rec}}$ is defined as $L_{\text{Rec}}=\lambda_{r}\cdot\omega\cdot L_{1}+(1-\lambda_{r})\cdot\omega\cdot L_{\text{SSIM}}$, where the $\lambda_{r}$ is a loss weight to balance the $L_{1}$ and $L_{D-SSIM}$. $\omega$ denotes the Binary Motion Mask (BMM) [27] that remove the distractor and marine snow artefacts. $L_{1}$ refers to the absolute error loss and $L_{\text{SSIM}}$ is the SSIM loss.
@@
The depth information plays a vital role as the depth $z$ information is involved in the estimation of the underwater medium parameters, as illustrated in Eq. 5. Thus, the $L_{\text{Depth}}$ is included and defined as, $L_{\text{Depth}}=\lambda_{d}\cdot\|\hat{D}-D\|+\lambda_{ca}\cdot(\sum_{ch\in\{r,g,b\}}\sum_{n\in\{d,b\}}\|z^{n}_{ch}-D\|)$, where $\lambda_{d}$ and $\lambda_{ca}$ are loss weights, $D$ is the pseudo depth map, obtained from Depth-Anything-V2 [31], and $\hat{D}$ is the rendered depth map from the RUSplatting. $z_{ch}^{n}$ denotes the inferred depth under direct ($d$) and backscatter ($b$) lighting conditions for each colour channel $ch$.
@@
Following [6], $L_{g}$ is built based on the assumption that in a natural scene, the average value of each colour channel of the image should be equal. With this restriction, it can promote the colour balance and aid in the accurate estimation of the medium parameters. $L_{g}=\sum_{ch\in\{r,g,b\}}\left\|\mu(J_{ch})-0.5\right\|_{2}^{2}$. Here, $J_{ch}$ denotes the restored pixel values for all three colour channels, ranging between 0 to 1. $\mu(\cdot)$ denotes the mean value of a channel. It is worth noting that the $0.5$ here corresponds to the ideal mean intensity for the colour channel under the grey-world assumption.

### v2

@@ Abstract
Reconstructing high-fidelity underwater scenes remains a challenging task due to light absorption, scattering, and limited visibility inherent in aquatic environments. This paper presents an enhanced Gaussian Splatting-based framework that improves both the visual quality and geometric accuracy of deep underwater rendering. We propose decoupled learning for RGB channels, guided by the physics of underwater attenuation, to enable more accurate colour restoration. To address sparse-view limitations and improve view consistency, we introduce a frame interpolation strategy with a novel adaptive weighting scheme. Additionally, we introduce a new loss function aimed at reducing noise while preserving edges, which is essential for deep-sea content. We also release a newly collected dataset, Submerged3D, captured specifically in deep-sea environments. Experimental results demonstrate that our framework consistently outperforms state-of-the-art methods with PSNR gains up to 1.90dB, delivering superior perceptual quality and robustness, and offering promising directions for marine robotics and underwater visual analytics. The code of RUSplatting is available at https://github.com/theflash987/RUSplatting and the dataset Submerged3D can be downloaded at https://zenodo.org/records/15482420.
@@ 1 Introduction
Advancements in Novel View Synthesis (NVS) have introduced new possibilities for reconstructing and visualising environments in 3D space from multi-view image data. These methods allow the generation of unseen viewpoints using limited training images. A notable breakthrough in NVS is Neural Radiance Fields (NeRF) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng], which implicitly models 3D scenes. Despite their impressive performance, NeRFs require substantial training time and are impractical for real-time or interactive applications [Chen and Wang(2024), Wu et al.(2024)Wu, Yuan, Zhang, Yang, Cao, Yan, and Gao]. Moreover, their representations are inherently difficult to interpret or modify intuitively [Wu et al.(2024)Wu, Yuan, Zhang, Yang, Cao, Yan, and Gao], limiting their flexibility in dynamic environments, such as underwater. In contrast, 3D Gaussian Splatting (3DGS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] reconstructs scenes explicitly using Gaussian point sets, offering faster training, higher rendering speeds and refresh rates, as well as more flexible scene manipulation.
@@
Nevertheless, most existing NeRF-based and 3DGS-based methods focus on reconstructing scenes in clear media and therefore struggle to produce high-quality results in hazy and distorted environments like underwater. Due to the physical properties of water, underwater imaging is affected by attenuation and scattering, where the former causes colour casting and the latter results in low contrast [Liu et al.(2024)Liu, Lu, Gu, Li, and Deng]. It is therefore essential to account for these properties when modelling the entire scene.
@@
In the face of these challenges, we propose a new framework: Robust Underwater scene representation using 3D Gaussian Splatting (RUSplatting), which demonstrates greater robustness than previous works, such as UW-GS [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull] and WaterSplatting [Li et al.(2025)Li, Song, Xu, Elsig, and Kulhanek]. Previous approaches integrate medium effects via affine transformations of Gaussian attributes but assume dense, temporally coherent input frames, limiting effectiveness under sparse-view conditions. They also do not address distortions common in deep-sea imaging, such as strong noise and severe colour imbalance.
@@
Our proposed RUSplatting framework, shown in Fig. 1, further advances underwater 3D reconstruction by explicitly decoupling the estimation of attenuation and scattering parameters for individual RGB channels instead of treating them uniformly, enabling more accurate and realistic colour restoration. As discussed in [Peng and Cosman(2017)], the light attenuation underwater has different degrees of changes due to the wavelength. Red light, which has a longer wavelength, is absorbed more rapidly than green and blue light. Furthermore, to mitigate image noise prevalent in deep-sea conditions, we introduce an Edge-Aware Smoothness Loss (ESL) to directly optimise the reconstruction pipeline, dynamically reducing noise while preserving critical geometric edges. Additionally, to address the inherent challenges of sparse-view datasets, we incorporate an Intermediate Frame Interpolation (IFI) mechanism, enriching viewpoint coverage and increasing initialised Gaussian points. However, since interpolated frames may contain artefacts that could deteriorate the learning process, we propose an Adaptive Frame Weighting (AFW) strategy that adjusts the loss based on the predictive uncertainty [Kendall and Gal(2017), Kendall et al.(2018)Kendall, Gal, and Cipolla]. Finally, we introduce a new dataset, Submerged3D, capturing detailed objects in deep-sea environments, supporting future research in underwater robotics, archaeology, and marine biology.
In summary, our key contributions are as follows:
@@
Extensive experiments on both existing and new datasets demonstrate that RUSplatting consistently outperforms state-of-the-art methods across various evaluation metrics, with PSNR improvements of up to 21.37%, 4.99% and 5.09% compared to SeeThru-NeRF [Levy et al.(2023)Levy, Peleg, Pearl, Rosenbaum, Akkaynak, Korman, and Treibitz], WaterSplatting [Li et al.(2025)Li, Song, Xu, Elsig, and Kulhanek] and UW-GS [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull], respectively.
@@ 2 Related Work
Processing underwater images remains a major challenge in underwater computer vision due to colour distortion and noise arising from water’s physical properties and its impact on light propagation [González-Sabbagh and Robles-Kelly(2023)]. The Underwater Image Formation Model (UIFM) [Chiang and Chen(2011), Drews et al.(2016)Drews, Nascimento, Botelho, and Campos] was introduced to describe this process, accounting for light absorption and backscattering. While the original UIFM assumes a homogeneous water column, the Revised UIFM (R-UIFM) [Akkaynak and Treibitz(2018)] improves realism by allowing spatial variability in water properties, making it widely adopted in underwater image processing. For example, PRWNet [Huo et al.(2021)Huo, Li, and Zhu] uses a deep learning-based multi-stage refinement approach that enhances image quality by addressing spatial and frequency domain degradations through Wavelet Boost Learning. However, the survey in [Huang et al.(2024)Huang, Wang, Seymour, Kovacs, Ellerbrock, Blackham, and Anantrasirichai] shows that applying enhancement prior to 3D reconstruction results in poorer performance compared to embedding the physical properties of water directly into the reconstruction process.
@@
While the NeRF-based and GS-based technologies have rapidly advanced, most struggle in scattering media such as fog or water. To tackle this limitation, ScatterNeRF [Ramazzina et al.(2023)Ramazzina, Bijelic, Walz, Sanvito, Scheuble, and Heide] was introduced to effectively separate objects from the scattering medium in foggy environments. SeaThru-NeRF [Levy et al.(2023)Levy, Peleg, Pearl, Rosenbaum, Akkaynak, Korman, and Treibitz] extends this line of research by estimating water medium coefficients using a novel architecture grounded in the Revised Underwater Image Formation Model (R-UIFM) [Akkaynak and Treibitz(2018)]. AquaNeRF [Gough et al.(2025)Gough, Azzarelli, Zhang, and Anantrasirichai] improves SeaThru-NeRF with single surface rendering to avoid distractors. $\bmvaOneDot[Tang et al$.(2024)Tang, Zhu, Wan, Xu, and Shi] propose a neural representation that incorporates dynamic rendering and tone mapping to enable more accurate reconstruction of underwater scenes under varying illumination conditions.
For 3DGS-based methods, UW-GS [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull] and WaterSplatting [Li et al.(2025)Li, Song, Xu, Elsig, and Kulhanek] employ MLP-based colour appearance modules to address colour casting. In contrast, SeaSplat [Yang et al.(2025)Yang, Leonard, and Girdhar] optimises backscatter parameters using a dark channel prior loss. UW-GS [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull], UDR-GS [Du et al.(2024)Du, Zhang, Zhang, Sun, and Lv], and Aquatic-GS [Liu et al.(2024)Liu, Lu, Gu, Li, and Deng] leverage DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] to estimate depth from 2D inputs and incorporate it into their loss functions. UW-GS [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull] further improves reconstruction of distant objects through a physics-driven density control module.
@@ 3 Preliminaries
The 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] employs a collection of spatial Gaussian cloud points to represent a 3D scene. To synthesise a novel view, these Gaussians are projected onto a 2D image plane. The projected Gaussian is defined as
@@
Compared to clear, onshore scenes, where scattering and absorption caused by the medium (i.e., air) can often be neglected, underwater scenes are heavily affected by these parameters, since they significantly degrade reconstruction quality [Tang et al.(2024)Tang, Zhu, Wan, Xu, and Shi]. Therefore, it is crucial to consider these medium parameters during reconstruction. As modelled in [Akkaynak and Treibitz(2018)], the intensity of the underwater image is defined as follows:
@@
Here $\beta^{d}$ and $\beta^{b}$ are two colour medium coefficients, and $z$ is the line-of-sight distance from the camera to the 3D scene point rather than the water depth.
@@ 4 Methodology
The detailed structure of the proposed framework is presented in Fig. 1. The model estimates underwater medium parameters using an MLP with encoded depth and viewpoint position information as inputs. The MLP $f_{\theta}(\gamma_{z}(z),\gamma_{\theta}(\mathbf{v}))$ predicts, for each Gaussian and channel, the parameters $\{{T^{D},T^{B},\beta^{d},\beta^{b},b}\}$, which are jointly optimised under the main loss function described in Section. 4.4. Accordingly, the colours of the spatial 3D Gaussians are modified through the affine transformation. Then, the modified Gaussians, including additional Gaussians introduced by the frame interpolation process with learnable uncertainty weights, are projected onto the 2D plane. During training, an edge-aware smoothness loss is integrated into the model as a part of the loss function to remove noise from underwater images, and in the meantime maintaining geometric details.
@@ 4.1 Colour-Channel Decoupling
As discussed in numerous underwater studies [Liu et al.(2024)Liu, Lu, Gu, Li, and Deng, Li and Anantrasirichai(2025), Huang et al.(2024)Huang, Wang, Seymour, Kovacs, Ellerbrock, Blackham, and Anantrasirichai], the attenuation rates differ across red, green, and blue light channels. Therefore, the medium parameters for the RGB channels cannot be treated identically. To overcome this and achieve more accurate colour reconstruction, we decouple the channel-wise learning of the underwater medium. Consequently, the attenuation ($T^{D}$) and backscatter ($T^{B}$) factors for each colour channel $ch\in\{r,g,b\}$ are defined as:
@@
Previous works use a 2-layer MLP [Lee et al.(2024)Lee, Rho, Sun, Ko, and Park] to estimate the complex underwater medium parameters, however, it may fall short in capturing such complicated patterns as the study indicated in [He et al.(2016)He, Zhang, Ren, and Sun]. Thus, we employ a deeper MLP with 5 layers, selected based on a hyperparameter sensitivity analysis (see supplementary material for details). The final corrected colour of Gaussian $i$, $c_{i,ch}^{m}(v)$, is obtained through an affine transformation, as presented below, where $c_{i,ch}(v)$ is the observed view-dependent colour.
@@
Here, $b_{i,ch}$ denotes a bias term specific to each Gaussian and channel, modelling the background light. It is assumed view-independent within the views observing Gaussian $i$, while view-dependent effects are captured by $c_{i,ch}(v)$. As Eq. 6 follows directly from Eq. 3, the observed channel-wise colour equals a scaled direct term plus an additive background light term, hence the mapping is affine.
@@ 4.2 Intermediate Frame Interpolation (IFI)
To address the limitations introduced by sparse views, we adopt a frame interpolation technique to synthesise intermediate frames between adjacent input images, leveraging them in two key aspects: i) enhancing the initial point cloud and ii) improving the RUSplatting optimisation, where we propose Adaptive Frame Weighting (AFW). Although various methods are available and should not significantly affect the final 3DGS results, we use Real-time Intermediate Flow Estimation (RIFE) [Huang et al.(2022)Huang, Zhang, Heng, Shi, and Zhou], a state-of-the-art deep learning-based approach that employs bidirectional optical flow and knowledge distillation to enable real-time processing.
@@
The initial point cloud (obtained using COLMAP [Schönberger and Frahm(2016)]) acts as a rough geometric framework for the final model. Its quality and density directly influence the accuracy of the rendering. By incorporating interpolated frames and re-running COLMAP on the augmented dataset, additional initialisation points are introduced, enabling a denser distribution of Gaussians and improving the overall reconstruction precision and detail. Without IFI, the model accesses an average of 17,635 initialisation points across all datasets. After adopting IFI, this number increases significantly by 55.7% to an average of 27,461 points. Detailed results for each dataset are provided in the supplementary material.
@@
In sparse view scenarios, adjacent frames are significantly spaced apart, making frame interpolation challenging. The large spatial shifts between frames may introduce artefacts in certain geometric structures of the interpolated frames, which may degrade the quality of 3D reconstruction during optimisation when these frames are used. Inspired by [Kendall and Gal(2017), Kendall et al.(2018)Kendall, Gal, and Cipolla], we propose an AFW mechanism, introducing an uncertainty learnable parameter $\gamma$ for each interpolated frame $f$ that is optimised along with the training. Accordingly, the loss function for the interpolated frame $L_{f}$ is updated as follows.
@@ 4.3 Edge-Aware Smoothness Loss (ESL)
where $(i,j)$ indexes the spatial coordinates in the image plane, $I$ denotes the input image and the $D$ denotes the pseudo depth map generated from the Depth-Anything-V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. The edge-aware weights $w^{x}_{i,j}$ and $w^{y}_{i,j}$, corresponding to the horizontal and vertical directions, respectively, are computed using depth discontinuities,
@@ 4.4 Loss Function
To evaluate the differences between the reconstructed image and the ground truth image, the $L_{\text{Rec}}$ is defined as $L_{\text{Rec}}=\lambda_{r}\cdot\omega\cdot L_{1}+(1-\lambda_{r})\cdot\omega\cdot L_{\text{SSIM}}$, where the $\lambda_{r}$ is a loss weight to balance the $L_{1}$ and $L_{D-SSIM}$. $\omega$ denotes the Binary Motion Mask (BMM) [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull] that remove the distractor and marine snow artefacts. $L_{1}$ refers to the absolute error loss and $L_{\text{SSIM}}$ is the SSIM loss.
@@
The depth information plays a vital role as the depth $z$ information is involved in the estimation of the underwater medium parameters, as illustrated in Eq. 5. Thus, the $L_{\text{Depth}}$ is included and defined as, $L_{\text{Depth}}=\lambda_{d}\cdot\|\hat{D}-D\|+\lambda_{ca}\cdot(\sum_{ch\in\{r,g,b\}}\sum_{n\in\{d,b\}}\|z^{n}_{ch}-D\|)$, where $\lambda_{d}$ and $\lambda_{ca}$ are loss weights, $D$ is the pseudo depth map, obtained from Depth-Anything-V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], and $\hat{D}$ is the rendered depth map from the RUSplatting. $z_{ch}^{n}$ denotes the inferred depth under direct ($d$) and backscatter ($b$) lighting conditions for each colour channel $ch$.
@@
Following [Fu et al.(2022)Fu, Lin, Yang, Chai, Sun, Huang, and Ding], $L_{g}$ is built based on the assumption that in a natural scene, the average value of each colour channel of the image should be equal. With this restriction, it can promote the colour balance and aid in the accurate estimation of the medium parameters. $L_{g}=\sum_{ch\in\{r,g,b\}}\left\|\mu(J_{ch})-0.5\right\|_{2}^{2}$. Here, $J_{ch}$ denotes the restored pixel values for all three colour channels, ranging between 0 to 1. $\mu(\cdot)$ denotes the mean value of a channel. It is worth noting that the $0.5$ here corresponds to the ideal mean intensity for the colour channel under the grey-world assumption.